OpenStudy (anonymous):

Please double check for me... For the function f(x) = (8-2x)2 ,find f-1 . Determine whether f-1 is a function

OpenStudy (anonymous):

I think the answer is going to be D??? Just not 100% sure

OpenStudy (phi):

yes

OpenStudy (anonymous):

y = (8-2x)^2 inv x = (8-2y)^2 sqrt(x) = 8-2y -sqrt(x) = -8+2y 8 -sqrt(x) = 2y (8 -sqrt(x))/2 = y becomes 8 (+/-) sqrt(x) ---------------- = y 2 since the graph of sqrt() looks like an elongated "C" a vertical line will pass thru it at more than one point, therefore, not a function so D

OpenStudy (anonymous):

hope that helps